OpenStudy (anonymous):

Hi everyone! I'm not sure why I can't remember how to do this but can someone help me take the derivative of sin^2(x) ? Thanks!

OpenStudy (anonymous):

\[\sin^2x=(\sin x)^2\] When using the chain rule, the "inside" function is sine, and the "outside" function is the squared. So, \[\frac{d}{dx}\sin^2x=2\sin x\cdot\left(\frac{d}{dx}\sin x\right)\]

OpenStudy (anonymous):

LOL...omg...I can't believe I forgot this! Thanks Giggles! :o)

OpenStudy (anonymous):

You're welcome!